OpenStudy (anonymous):

given that x2+ax+b=(x-7)2-a work out a and b

OpenStudy (bahrom7893):

x^2 + ax + b = 0x^2 + 2x - 14 - a

OpenStudy (bahrom7893):

x^2 + ax + b = 0x^2 + 2x + ( - 14 - a )

jhonyy9 (jhonyy9):

(x-7)^2 = x^2 -14x +49 - so than can you continue it ?

OpenStudy (bahrom7893):

ohhhh lol sorry johnny is right.. guys please use the correct notation haha

OpenStudy (bahrom7893):

I assumed it was ()*2

OpenStudy (anonymous):

Can u explain the steps please?

OpenStudy (bahrom7893):

x2+ax+b=(x-7)2-a x^2 + ax + b = x^2 - 14x + 49 - a, right?

OpenStudy (bahrom7893):

Subtract x^2 from both sides: ax+b = -14x + (49-a)

OpenStudy (bahrom7893):